OpenStudy (anonymous):

Please help me out. Which pair of functions is not a pair of inverse functions?

OpenStudy (anonymous):

f(x)=x+1/6 and g(x)=6x−1 f(x)=x−4/19 and g(x)=19x+4 \[f(x)=x^5 and g(x)=\sqrt[5]{x}\] f(x)=x/x+20 and g(x)=20xx−1

OpenStudy (anonymous):

the last one is g(x)=20x/x-1

OpenStudy (anonymous):

@thomas5267

OpenStudy (anonymous):

@nincompoop